OpenStudy (anonymous):

A block accelerates at 3.7 m/s2 down a plane inclined at an angle 27.0. Find μk between the block and the in- clined plane. The acceleration of gravity is 9.81 m/s2 .

OpenStudy (anonymous):

Ok once again worked out.... Tcos32=6.5 * 9.8 T=75.12 75.12sin32=6.5 * a a=6.12 Please tell me if this is the right answer....

OpenStudy (anonymous):

ok sorry that is def the wrong question this is the RIGHT question....sorry guys..

OpenStudy (anonymous):

A 6.5 kg object hangs at one end of a rope that is attached to a support on a railroad boxcar. When the car accelerates to the right, the rope makes an angle of 32 with the vertical The acceleration of gravity is 9.8 m/s2 .Find the acceleration of the car. (Hint: ~aobject = ~acar) Answer in units of m/s2

OpenStudy (anonymous):

and actually i dont even have an idea on how to do the question I originally posted.

OpenStudy (anonymous):

Your work looks good. Let me analyze the original question.

OpenStudy (anonymous):

Find the gravitational force acting along the inclined plane and the gravitation force acting normal to the inclined plane. The balance of the force acting along the inclined plane and the friction force will balance to equal the acceleration the object is experiencing. \[m \cdot a = F_{ip} - F_f\]

OpenStudy (anonymous):

ok normal force = m*g cos theta....but how do you find mass???

OpenStudy (anonymous):

Force * Acceleration ahhh duh idiot*

OpenStudy (anonymous):

but then that would be a unit-less number...that cant be right... oh come back my savior

OpenStudy (anonymous):

so i tried this : m * a = m*g*sinA- uk*m*g*cosA a= g*sinA - uk *g*cosA 3.7=9.81 * sin 27 -uk * 9.81 * cos 27 3.7 = 4.45 - 8.74uk 3.7 - 4.45 = -8.74uk .09 = uk ???? yes no maybe??

OpenStudy (anonymous):

Looks good.

OpenStudy (anonymous):

wrong ?????? help.